How long does it take to travel .2 miles at 35 miles per hour?
Found 2 solutions by rfer, Alan3354:
Answer by rfer(12654)

.2/35=0.0057 hrs

Answer by Alan3354(30993)

d = rt
t = d/r = 0.2/35
t =~ 0.005714 hours